OpenStudy (yumyum247):

so i have a quick question...while taking the derivative of an exponential function, where f(x)= 2^x. the derivative f'(x)=is k(fx) or f'(x)=k(2^x). my question is what's k and its purpose??

OpenStudy (irishboy123):

you need to switch into natural logs and e before you can play with the usual calculus rules \(y = 2^x\) \(ln y = x ln 2\) off you go

OpenStudy (sshayer):

\[\frac{ d }{ dx }(a^x)=a^x \ln a\]

OpenStudy (kainui):

I think your notation is off, because this doesn't seem to be meaningful, " f'(x)=is k(fx) or f'(x)=k(2^x)." But if your function is \(f(x)=2^x\) then maybe they are suggesting you put in a constant \(k\) in so that you get: \(f(kx)=2^{kx}\). Then you can use the chain rule to differentiate this but I need more info or something cause it's sorta ambiguous what you're doing.

OpenStudy (irishboy123):

it's purpose is the log switch to logs that work for calculus see @sshayer 's thing above